OpenStudy (anonymous):

A disc jockey must play 12 commercial spots during 1 hour of a radio show. Each commercial is either 30 seconds, or 60 seconds long. If the total time is 1 hour 8 min, how many 30 and sixty seconds commercials were played?

OpenStudy (anonymous):

We will denote 30 second commercial as 'a' 60 second commercial as 'b'

OpenStudy (anonymous):

There are 12 commercial spot a+b=12

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you check all your numbers again?

OpenStudy (anonymous):

if the total commercial time during the 1 hour is 8 minutes

OpenStudy (anonymous):

i wrote it wrong

OpenStudy (anonymous):

\[a+b\text{=}12\] \[\frac{1}{120}a+\frac{1}{60}b=\frac{8}{60}\] Solve for a , b a=8 b=4